Hydrochloric acid is sold as a concentrated aqueous solution. If the concentration of commercial is 12M and its density is 1.18g/cm3 , calculate the following: (a) the molality of the HCl solution (b) the weight percent of HCl in the solution.

Please show the work on how to arrive at these answers -

a. 16.2m

b. 37.1%
Answer #1

12.0 M means 12.0 moles in 1 L of solution

Mass HCl = 12.0 mol x 36.46 g/mol =437.5 g

Mass of 1 L of solution = 1180 g

Mass water = 1180 - 437.5 = 742.5 g = 0.7425 Kg

Now, a) Molality of HCl = 12.0 / 0.7425 = 16.2 m

b) weight percent of HCl = 437.5 x 100 / 1180 =37.07=37.1 %